Solve logx+log(x-3)=log108
Answer by LinnW(1048)

log(x)+log(x-3)=log(108)
log(x*(x-3)) = log(108)
x*(x-3) = 108
x^2 - 3x = 108
x^2 - 3x - 108 = 0
(x - 12)(x + 9) = 0
x = 12 or x = -9